A stone of weight 32g falls from the top of a 250m high cliff.i.Calculate the work done by the force of gravity in pulling the stone to the foot of the cliff. 


1
Expert's answer
2021-04-18T19:32:12-0400

The work is scalar product of F and displacement. Since the stone displacement is in the same direction with the force, scalar product becomes just an ordinary product.

"A = F \cdot l = mgl =32 \cdot 10^{-3} \cdot 9.81 \cdot 250 = 78.48 \, J"


Answer: 78.5 J